# Dynamic Scheduling for Vehicle-to-Vehicle Communications Enhanced Federated Learning

### v1

@@ Abstract
Leveraging the computing and sensing capabilities of vehicles, vehicular federated learning (VFL) has been applied to edge training for connected vehicles. The dynamic and interconnected nature of vehicular networks presents unique opportunities to harness direct vehicle-to-vehicle (V2V) communications, enhancing VFL training efficiency. In this paper, we formulate a stochastic optimization problem to optimize the VFL training performance, considering the energy constraints and mobility of vehicles, and propose a V2V-enhanced dynamic scheduling (VEDS) algorithm to solve it. The model aggregation requirements of VFL and the limited transmission time due to mobility result in a stepwise objective function, which presents challenges in solving the problem. We thus propose a derivative-based drift-plus-penalty method to convert the long-term stochastic optimization problem to an online mixed integer nonlinear programming (MINLP) problem, and provide a theoretical analysis to bound the performance gap between the online solution and the offline optimal solution. Further analysis of the scheduling priority reduces the original problem into a set of convex optimization problems, which are efficiently solved using the interior-point method. Experimental results demonstrate that compared with the state-of-the-art benchmarks, the proposed algorithm enhances the image classification accuracy on the CIFAR-10 dataset by $3.18\%$ and reduces the average displacement errors on the Argoverse trajectory prediction dataset by $10.21\%$.
@@ II Related Works
Many studies have explored the application of federated learning (FL) in wireless networks [20], addressing critical issues such as wireless resource management[21, 22, 23, 24, 25, 26, 27], compression and sparsification[28, 29, 30, 31, 32], and training algorithm design [33, 34, 35]. However, these studies rarely consider the unique characteristics of vehicular networks, such as high mobility and rapidly changing channel conditions.
More recent studies have begun to investigate FL in vehicular networks. These studies recognize the challenges posed by the high mobility of vehicles and the dynamic nature of vehicular environments [36, 5, 7, 37]. In [36], the impact of vehicle mobility on data quality, such as noise, motion blur, and distortion, is considered, and a resource optimization and vehicle selection scheme is proposed in the context of VFL. The proposed scheme dynamically schedules vehicles with higher image quality, increasing the convergence rate and reducing the time and energy consumption in FL training. In [5], the short-lived connections between vehicles and RSUs are considered, and a mobility-aware optimization algorithm is proposed. The proposed algorithm enhances the convergence performance of VFL by optimizing the duration of each training round and the number of local iterations. In [7, 37], the impact of rapidly time-varying channels resulting from vehicle mobility is considered. Specifically, a mobility and channel dynamic aware FL (MADCA-FL) scheme is proposed in [7], which optimizes the success probability of vehicle selection and model parameter updating based on the analysis of vehicle mobility and channel dynamics. In [37], a more realistic scenario is explored within a 5G new radio framework, and a joint VFL and radio access technology parameter optimization scheme is proposed under the constraints of delay, energy, and cost, aiming to maximize the successful transmission rate of locally trained models. However, most existing studies focus on V2I aggregation, overlooking the potential of harnessing V2V sidelinks to enhance the VFL training efficiency.
@@ III-A VFL Model
We consider a VFL system as shown in Fig. 1, where an RSU (indexed by $r$ in the following) orchestrates the training of a neural network model $\boldsymbol{w}$ with the assistance of vehicles that enter its coverage area. During the $k^{\text{th}}$ training round, the vehicles that possess local datasets and are willing to participate in the collaborative training of the neural network model are referred to as SOVs, denoted by $\mathcal{S}_{k}$. The vehicles that do not participate in model training, but have communication capabilities and can help SOVs upload the models are referred to as OPVs, denoted by $\mathcal{U}_{k}$.
Each SOV $m\in\mathcal{S}_{k}$ holds a local dataset with an associated distribution $\mathcal{D}_{m}$ over the space of samples $\mathcal{X}_{m}$. For each data sample $\boldsymbol{x}\in\mathcal{X}_{m}$, a loss function $f(\boldsymbol{w};\boldsymbol{x})$ is used to measure the fitting performance of the model vector $\boldsymbol{w}$. The local loss function of vehicle $m$ is defined as the average loss over the distribution $\mathcal{D}_{m}$, i.e., $f_{m}(\boldsymbol{w})\triangleq\underset{\boldsymbol{x}\sim\mathcal{D}_{m}}{\operatorname*{\mathbb{E}}}[f(\boldsymbol{w};\boldsymbol{x})].$
@@ III-A1 Local Updates
$\boldsymbol{w}_{m,k}=\boldsymbol{w}_{k-1}-\frac{\eta_{k}}{B_{k}}\sum_{\boldsymbol{x}\in\mathcal{B}_{m,k}}\nabla f\left(\boldsymbol{w}_{k-1};\boldsymbol{x}\right),$ (2)
where $\eta_{k}$ is the learning rate, $\mathcal{B}_{m,k}$ is a subset randomly sampled from the sample space $\mathcal{X}_{m}$. We assume that the batch size of all SOVs is the same, and denote it by $B_{k}=|\mathcal{B}_{m,k}|$.
@@ III-A3 Model Aggregation
$\boldsymbol{w}_{k}=\frac{\sum_{m\in\hat{\mathcal{S}}_{k}}|\mathcal{D}_{m}|\boldsymbol{w}_{m,k}}{\sum_{m\in\hat{\mathcal{S}}_{k}}|\mathcal{D}_{m}|},$ (3)
@@ III-B Computation Model
We adopt a standard computation model [38] [39] for local updates. The total workload for computing local updates for each vehicle is $N_{\text{flop}}B_{k}$, where $N_{\text{flop}}$ is the number of floating point operations (FLOPs) needed for processing each sample. Further, we define $l_{m,k}$ (in cycle/s) as the clock frequency of the vehicular processor in round $k$. Hence, the computation latency for updating the local model is determined as follows:
$t^{\text{cp}}_{m,k}=\frac{N_{\text{flop}}B_{k}}{l_{m,k}},$
@@
$e^{\text{cp}}_{m,k}=\rho l_{m,k}^{2}N_{\text{flop}}B_{k},$
@@ III-C Communication Model
$R_{m}^{\text{DT}}(t)=\beta\log_{2}\left(1+\frac{p_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}\right),$
@@
For COT, the scheduled SOV uses the first half of the slot to transmit its model parameters to the OPVs, and the OPVs use distributed space-time code (DSTC) [40] to relay the model parameters to the RSU in the second half of the slot, as shown in Fig. 2. We use $\boldsymbol{u}(t)=[u_{1}(t),...,u_{|\mathcal{U}_{k}|}(t)]$ to denote the OPV scheduling decision in slot $t$, where $u_{n}(t)=1$ if the OPV $n\in\mathcal{U}_{k}$ is scheduled for COT, and $u_{n}(t)=0$ otherwise. $u_{n}(t)$ has the binary constraint:
@@
$R_{m}^{\text{COT}}(t)=\beta\log_{2}\bigg{(}1+\frac{p_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}$
$+\sum_{n\in\mathcal{U}_{k}}\frac{u_{n}(t)p_{n}(t)|h_{n,r}(t)|^{2}}{\beta N_{0}}\bigg{)}.$
@@
$R_{m,n}^{\text{COT-V}}(t)=\beta\log_{2}\bigg{(}1+\frac{p_{m}(t)|h_{m,n}(t)|^{2}}{\beta N_{0}}\bigg{)},$
@@
$\boldsymbol{w}_{k}=\frac{\sum_{m\in{\mathcal{S}}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)|\mathcal{D}_{m}|\boldsymbol{w}_{m,k}}{\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)|\mathcal{D}_{m}|}.$ (11)
@@ IV-A Convergence Analysis
Assumption 1: The local loss function $f_{m}(\boldsymbol{w})$ is $L$-smooth for each SOV $m\in\mathcal{S}_{k}$ in each round $k\in\mathcal{K}$, i.e.,
@@
Assumption 2: The local loss function $f_{m}(\boldsymbol{w})$ is $\mu$-strongly convex for each SOV $m\in\mathcal{S}_{k}$ in each round $k\in\mathcal{K}$, i.e.,
$f_{m}(\boldsymbol{w}_{k}$ $)-f_{m}(\boldsymbol{w}_{k-1})$
Assumption 3: The stochastic gradient is unbiased and variance-bounded, i.e.,
$\underset{\boldsymbol{x}\sim\mathcal{D}_{m}}{\operatorname*{\mathbb{E}}}\left[\nabla f(\boldsymbol{w};\boldsymbol{x})\right]=\underset{m\sim\mathcal{P}_{m}}{\operatorname*{\mathbb{E}}}\left[\nabla f_{m}(\boldsymbol{w})\right]=\nabla F(\boldsymbol{w}),$
$\underset{\boldsymbol{x}\sim\mathcal{D}_{m}}{\operatorname*{\mathbb{E}}}\left[\left\|\nabla f(\boldsymbol{w};\boldsymbol{x})-\nabla F(\boldsymbol{w})\right\|^{2}\right]\leq G^{2}.$
@@
Theorem 1. After $K$ round of training, the difference between $F({\boldsymbol{w}}_{K})$ and the optimal global loss function $F(\boldsymbol{w}^{*})$ is upper bounded by
@@ IV-B Problem Formulation
Based on Theorem 1, we alternatively minimize the upper bound of $\mathbb{E}[F({\boldsymbol{w}}_{K})]-F(\boldsymbol{w}^{*})$ in (13), which is equivalent to minimizing $\frac{\eta_{k}}{2}\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}$ in each round $k\in\mathcal{K}$. The optimization problem is formulated as
$P0:$ $\underset{\boldsymbol{S}_{k},\boldsymbol{c}_{k},\boldsymbol{U}_{k},\boldsymbol{P}_{k}}{\min}\ \frac{\eta_{k}}{2}\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}$ (14a)
@@
Since $\eta_{k}$, $G$ and $B_{k}$ are all constants in round $k$, minimizing (14a) is equivalent to minimizing $\frac{1}{\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}$. Also, since $\frac{1}{\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}>0$, there is
@@ V-A Transformation of the stochastic optimization problem
$q_{m}^{\text{SOV}}(t+1)=\max\left\{q_{m}^{\text{SOV}}(t)+e_{m}^{\text{cm}}(t)-\frac{E_{m}^{\text{cons}}-e_{m,k}^{\text{cp}}}{T},0\right\}.$ (19)
@@
$q_{n}^{\text{OPV}}(t+1)=\max\left\{q_{n}^{\text{OPV}}(t)+e_{n}^{\text{cm}}(t)-\frac{E_{n}^{\text{cons}}}{T},0\right\}.$ (20)
@@
$\quad\quad\quad-\sum_{m\in\mathcal{S}_{k}}q_{m}^{\text{SOV}}e^{\text{cm}}_{m}(t)-\sum_{n\in\mathcal{U}_{k}}q_{n}^{\text{OPV}}e^{\text{cm}}_{n}(t)$ (21a)
@@
Proof: See Appendix C. $\square$
@@
V-B Direct Transmission Problem
s.t. $0\leq p_{m}(t)\leq p^{\text{max}}_{m}.$ (25b)
$p_{m}^{*}(t)=\left[\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\beta}{q_{m}^{\text{SOV}}(t)}-\frac{\beta N_{0}}{|h_{m,r}(t)|^{2}}\right]^{p^{\text{max}}_{m}}_{0},$ (26)
Proof: See Appendix D. $\square$
@@ V-B Direct Transmission Problem
           Solving $P4$ based to obtain $\boldsymbol{p}(t)$;

### v2

@@ Abstract
Leveraging the computing and sensing capabilities of vehicles, vehicular federated learning (VFL) has been applied to edge training for connected vehicles. The dynamic and interconnected nature of vehicular networks presents unique opportunities to harness direct vehicle-to-vehicle (V2V) communications, enhancing VFL training efficiency. In this paper, we formulate a stochastic optimization problem to optimize the VFL training performance, considering the energy constraints and mobility of vehicles, and propose a V2V-enhanced dynamic scheduling (VEDS) algorithm to solve it. The model aggregation requirements of VFL and the limited transmission time due to mobility result in a stepwise objective function, which presents challenges in solving the problem. We thus propose a derivative-based drift-plus-penalty method to convert the long-term stochastic optimization problem to an online mixed integer nonlinear programming (MINLP) problem, and provide a theoretical analysis to bound the performance gap between the online solution and the offline optimal solution. Further analysis of the scheduling priority reduces the original problem into a set of convex optimization problems, which are efficiently solved using the interior-point method. Experimental results demonstrate that compared with the state-of-the-art benchmarks, the proposed algorithm enhances the image classification accuracy on the CIFAR-10 dataset by $4.20\%$ and reduces the average displacement errors on the Argoverse trajectory prediction dataset by $9.82\%$.
@@ II Related Works
Many studies have explored the application of federated learning (FL) in wireless networks [20], addressing critical issues such as wireless resource management[21, 22, 23, 24, 25, 26, 27, 28], compression and sparsification[29, 30, 31, 32, 33], and training algorithm design [34, 35, 36, 37, 38]. However, these studies rarely consider the unique characteristics of vehicular networks, such as high mobility and rapidly changing channel conditions.
More recent studies have begun to investigate FL in vehicular networks. These studies recognize the challenges posed by the high mobility of vehicles and the dynamic nature of vehicular environments [39, 5, 6, 40, 7, 41]. In [39], the impact of vehicle mobility on data quality, such as noise, motion blur, and distortion, is considered, and a resource optimization and vehicle selection scheme is proposed in the context of VFL. The proposed scheme dynamically schedules vehicles with higher image quality, increasing the convergence rate and reducing the time and energy consumption in FL training. In [5], the short-lived connections between vehicles and RSUs are considered, and a mobility-aware optimization algorithm is proposed. The proposed algorithm enhances the convergence performance of VFL by optimizing the duration of each training round and the number of local iterations. In [6], a convergence analysis is provided to evaluate the impact of mobile users leaving the coverage area. A mobility-aware solution is proposed to mitigate the impact of mobility by clustering users with similar mobility patterns, which helps stabilize the model updates. Ref. [40] extends the analysis to the scenarios with heterogeneous data distributions, showing that mobility exacerbates the divergence of data heterogeneity, and therefore degrades the FL convergence. In [7, 41], the impact of rapidly time-varying channels resulting from vehicle mobility is considered. Specifically, a mobility and channel dynamic aware FL (MADCA-FL) scheme is proposed in [7], which optimizes the success probability of vehicle selection and model parameter updating based on the analysis of vehicle mobility and channel dynamics. In [41], a more realistic scenario is explored within a 5G new radio framework, and a joint VFL and radio access technology parameter optimization scheme is proposed under the constraints of delay, energy, and cost, aiming to maximize the successful transmission rate of locally trained models. However, most existing studies focus on V2I aggregation, overlooking the potential of harnessing V2V sidelinks to enhance the VFL training efficiency.
@@ III-A VFL Model
We consider a VFL system as shown in Fig. 1, where an RSU (indexed by $r$ in the following) orchestrates the training of a neural network model $\boldsymbol{w}$ with the assistance of vehicles that enter its coverage area. During the $k^{\text{th}}$ training round, the vehicles that possess local datasets and are willing to participate in the collaborative training of the neural network model are referred to as SOVs, denoted by $\mathcal{S}_{k}$. The vehicles that do not participate in model training, but have communication capabilities and can help SOVs upload the models are referred to as OPVs, denoted by $\mathcal{U}_{k}$. The RSU broadcasts control information, including resource allocation and scheduling decisions, while vehicles perform tasks like model uploading or relaying based on these instructions. Vehicles also report status information to the RSU, such as transmitted model parameters, channel conditions, and remaining battery levels.
Each SOV $m\in\mathcal{S}_{k}$ holds a local dataset with an associated distribution $\mathcal{D}_{m}$ over the space of samples $\mathcal{X}_{m}$. For each data sample $\boldsymbol{x}\in\mathcal{X}_{m}$, a loss function $f(\boldsymbol{w},\boldsymbol{x})$ is used to measure the fitting performance of the model vector $\boldsymbol{w}$. The local loss function of vehicle $m$ is defined as the average loss over the distribution $\mathcal{D}_{m}$, i.e., $f_{m}(\boldsymbol{w})\triangleq\underset{\boldsymbol{x}\sim\mathcal{D}_{m}}{\operatorname*{\mathbb{E}}}[f(\boldsymbol{w},\boldsymbol{x})].$
@@ III-A1 Local Updates
$\boldsymbol{w}_{m,k}=\boldsymbol{w}_{k-1}-\frac{\eta_{k}}{B_{m,k}}\sum_{\boldsymbol{x}\in\mathcal{B}_{m,k}}\nabla f\left(\boldsymbol{w}_{k-1},\boldsymbol{x}\right),$ (2)
where $\eta_{k}$ is the learning rate, $\mathcal{B}_{m,k}$ is a minibatch randomly sampled from the sample space $\mathcal{X}_{m}$. We assume that the batch size of all SOVs is the same, and denote it by $B_{k}=|\mathcal{B}_{m,k}|$.
@@ III-A3 Model Aggregation
$\boldsymbol{w}_{k}=\frac{1}{|\hat{\mathcal{S}}_{k}|}\sum_{m\in\hat{\mathcal{S}}_{k}}\boldsymbol{w}_{m,k},$ (3)
@@ III-B Computation Model
We adopt a standard computation model [42] [43] for local updates. The total workload for computing local updates for each vehicle is $N_{\text{flop}}B_{m,k}$, where $N_{\text{flop}}$ is the number of floating point operations (FLOPs) needed for processing each sample. Further, we define $l_{m,k}$ (in cycle/s) as the clock frequency of the vehicular processor in round $k$. Hence, the computation latency for updating the local model is determined as follows:
$t^{\text{cp}}_{m,k}=\frac{N_{\text{flop}}B_{m,k}}{l_{m,k}},$
@@
$e^{\text{cp}}_{m,k}=\rho l_{m,k}^{2}N_{\text{flop}}B_{m,k},$
@@ III-C Communication Model
$R_{m}^{\text{DT}}(t)=\beta\log_{2}\left(1+\frac{p_{m}(t)h_{m,r}(t)}{\beta N_{0}}\right),$
@@
For COT, the scheduled SOV uses the first half of the slot to broadcast its model parameters to the OPVs and the RSU. The OPVs use distributed space-time code (DSTC) [44, 45, 46] to relay the model parameters to the RSU in the second half of the slot, as shown in Fig. 2. DSTC is a cooperative communication technique that exploits spatial diversity to enhance transmission reliability by enabling multiple single-antenna OPVs to form a virtual multi-antenna system and improving robustness against channel fading.
We use $\boldsymbol{u}(t)=[u_{1}(t),...,u_{|\mathcal{U}_{k}|}(t)]$ to denote the OPV scheduling decision in slot $t$, where $u_{n}(t)=1$ if the OPV $n\in\mathcal{U}_{k}$ is scheduled for COT, and $u_{n}(t)=0$ otherwise. $u_{n}(t)$ has the binary constraint:
@@
$R_{m}^{\text{COT}}(t)=\beta\log_{2}\bigg{(}1+\frac{p_{m}(t)h_{m,r}(t)}{\beta N_{0}}$
$+\sum_{n\in\mathcal{U}_{k}}\frac{u_{n}(t)p_{n}(t)h_{n,r}(t)}{\beta N_{0}}\bigg{)}.$
@@
$R_{m,n}^{\text{COT-V}}(t)=\beta\log_{2}\bigg{(}1+\frac{p_{m}(t)h_{m,n}(t)}{\beta N_{0}}\bigg{)},$
@@
$\boldsymbol{w}_{k}=\frac{\sum_{m\in{\mathcal{S}}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)\boldsymbol{w}_{m,k}}{\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}.$ (11)
@@ IV-A Convergence Analysis
Assumption 1: The local loss function $f_{m}(\boldsymbol{w})$ of each SOV $m$ is $L$-smooth, i.e.,
@@
Assumption 2: The stochastic gradient of each SOV $m$ is variance-bounded, i.e.,
$\underset{\boldsymbol{x}\sim\mathcal{D}_{m}}{\operatorname*{\mathbb{E}}}\left[\left\|\nabla f(\boldsymbol{w},\boldsymbol{x})-\nabla f_{m}(\boldsymbol{w})\right\|^{2}\right]\leq G^{2}.$
@@
We derive the convergence bound for the convex case, assuming that the loss function of each SOV $m$ is $\mu$-strongly convex, i.e.,
$f_{m}(\boldsymbol{w}_{k}$ $)-f_{m}(\boldsymbol{w}_{k-1})$
@@
Theorem 1. (Convex) After $K$ round of training, the difference between $F({\boldsymbol{w}}_{K})$ and the optimal global loss function $F(\boldsymbol{w}^{*})$ is upper bounded by
@@
Further, we extend the convergence analysis to cases where the loss function is non-convex, which commonly occurs in ML models such as neural networks. Setting $\eta_{k}=\eta$ and based on lemma 1, we have the following theorem.
Theorem 2. (Non-convex) After $K$ rounds of training, the expected gradient norm of the global loss function is upper bounded by
$\frac{1}{K}$ $\sum_{k=1}^{K}\mathbb{E}\left\|\nabla F(\boldsymbol{w}_{k-1})\right\|^{2}\leq\frac{2L\left(F(\boldsymbol{w}_{0})-F(\boldsymbol{w}^{*})\right)}{K^{1/2}}$ (14)
$+\frac{1}{K^{3/2}}\sum_{k=1}^{K}\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}.$
Proof: See Appendix C. $\square$
Theorems 1 and 2 show that the convergence bound of VFL improves as the number of successful uploads per round, $\sum_{m\in\mathcal{S}k}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)$, increases. Our algorithm is designed to maximize the number of successful uploads in each round. The number of SOVs in each round $|\mathcal{S}_{k}|$ in each round serves as an upper limit on successful uploads, which in turn impacts the algorithm performance by constraining the maximum number of successful uploads.
@@ IV-B Problem Formulation
Based on the theorems above, we alternatively minimize the upper bound of $\mathbb{E}[F({\boldsymbol{w}}_{K})]-F(\boldsymbol{w}^{*})$ in (13) for convex loss functions, and the upper bound of $\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left\|\nabla F(\boldsymbol{w}_{k-1})\right\|^{2}$ in (14) for non-convex loss functions. In both cases, this is equivalent to minimizing $\frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}$ in each round $k\in\mathcal{K}$. The optimization problem is formulated as
$P0:$ $\underset{\boldsymbol{S}_{k},\boldsymbol{c}_{k},\boldsymbol{U}_{k},\boldsymbol{P}_{k}}{\min}\ \frac{G^{2}}{B_{k}\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}$ (15a)
@@
Since $G$ and $B_{k}$ are constants, minimizing (15a) is equivalent to minimizing $\frac{1}{\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}$. Also, since $\frac{1}{\sum_{m\in\mathcal{S}_{k}}\mathbb{I}\left(\sum_{t\in\mathcal{T}_{k}}z_{m}(t)\geq Q\right)}>0$, there is
@@ V-A Transformation of the stochastic optimization problem
$q_{m}^{\text{SOV}}(t+1)=\max\left\{q_{m}^{\text{SOV}}(t)+e_{m}^{\text{cm}}(t)-\frac{E_{m}^{\text{cons}}-e_{m,k}^{\text{cp}}}{T_{k}},0\right\}.$ (20)
@@
$q_{n}^{\text{OPV}}(t+1)=\max\left\{q_{n}^{\text{OPV}}(t)+e_{n}^{\text{cm}}(t)-\frac{E_{n}^{\text{cons}}}{T_{k}},0\right\}.$ (21)
@@
$\quad\quad-\sum_{m\in\mathcal{S}_{k}}q_{m}^{\text{SOV}}(t)e^{\text{cm}}_{m}(t)-\sum_{n\in\mathcal{U}_{k}}q_{n}^{\text{OPV}}(t)e^{\text{cm}}_{n}(t)$ (22a)
@@
Proof: See Appendix D. $\square$
@@
We remark here that, the scheduling and resource allocation decisions are mainly based on the amount of model transmitted, channel state, and virtual queues. As the number of vehicles increases or the VFL model size grows, limited communication and energy resources may result in incomplete model transmissions. The derivative-based drift-plus-penalty method prioritizes scheduling vehicles with better channel conditions and those that have already uploaded a large portion of their models. By focusing resources on these vehicles, the algorithm ensures that they can complete their uploads successfully, thereby maintaining overall performance under constrained resources.
@@
           Solve $P4$ to obtain $\boldsymbol{p}(t)$;
@@
V-B Direct Transmission Problem
s.t. $0\leq p_{m}(t)\leq p^{\text{max}}_{m}.$ (26b)
$p_{m}^{*}(t)=\left[\frac{V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\beta}{q_{m}^{\text{SOV}}(t)}-\frac{\beta N_{0}}{h_{m,r}(t)}\right]^{p^{\text{max}}_{m}}_{0},$ (27)
Proof: See Appendix E. $\square$